OpenStudy (anonymous):

passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to ?

OpenStudy (anonymous):

A)5x – 3y = 0 B)-x + 3y = 0 C)-5x – 3y = 0 D)3x + 5y = 0 E)-3x + 5y = 0

OpenStudy (phi):

I would find the slope between points A and B that is, find change in y and divide that by change in x